### v1

# Exploring decay and -delayed neutron emission in exotic 46,47Cl isotopes

Abstract
In this paper, and -delayed neutron decays of 46,47Cl are reported from an experiment carried out at the National Superconducting Cyclotron Laboratory using the Beta Counting System. The half-lives of both 46Cl and 47Cl were extracted. Based on the delayed -ray transitions observed, the level structure of 46Ar was determined. Completely different sets of excited states above the first state in 46Ar were populated in the 46Cl and 47Cl decay channels. Two new -ray transitions in 47Ar were identified from the very weak 47Cl decay. Furthermore, 46Cl and 47Cl were also observed to yield different population patterns for levels in 45Ar, including states of different parities. The experimental results allow us to address some of the open questions related to the delayed neutron emission process. For isotopes with large neutron excess and high values, delayed neutron emission remains an important decay mode and can be utilized as a powerful spectroscopic tool. Experimental results were compared with shell-model calculations using the FSU and effective interactions.
I Introduction
Experimental investigations away from the valley of stability have highlighted unique and interesting properties of exotic nuclei, including neutron halos, neutron skins, and the disappearance of the standard magic numbers. For extremely neutron-rich nuclei, -delayed neutron emission becomes an important pathway in the decay process. With increasing neutron number, the -value for decay increases and, simultaneously, the neutron-separation energy () in the daughter nucleus decreases. This leads to an increase in the population of neutron-unbound states in the allowed decay process. Additionally, for nuclei with a large excess of neutrons, the parent and daughter nuclei often have different ground-state parities due to the valence protons and neutrons occupying opposite-parity shells. As a result, 1-particle, 1-hole () states, which typically lie at high energies, are preferentially populated in allowed decay. This, in part, drives the shifting of the Gamow-Teller strength above the neutron-emission threshold. These states typically decay by emitting a neutron, leading to a daughter nucleus with , though, in principle, decay by -ray emission cannot be completely ruled out. For extremely neutron-rich nuclei, some Gamow-Teller strength can be located above the two-neutron separation energy, which will result in the emission of more than one neutron following decay.
The -delayed neutron emission process is often modeled in two steps, where the neutron-rich precursor nucleus undergoes decay to a highly-excited state in the daughter nucleus which then subsequently emits a neutron. It is assumed that neutron emission occurs from an equilibrated system completely independent of the formation process. Therefore, it is a statistical process and depends only on the spin, parity, and excitation energy of the level in the daughter nucleus. With very limited experimental information for confirmation, this assumption may be too simplistic, especially for lighter nuclei, such as those in the shell where the density of states above is relatively low, putting the statistical decay assumption in doubt [1]. For example, there is some experimental evidence that points toward neutron emission proceeding from an intermediate doorway state with a large probability for neutron decay [2]. The quantum dynamics of sequential decay proceeding via unbound neutron resonances is of significant theoretical interest because the intermediate state partially retains the memory of its initial creation. This evolving state, which comprises both resonant and background components, is non-stationary and therefore subject to internal mixing as well as explicit decay [3, 4].
Thus, -delayed neutron emission remains an important and unresolved challenge for both experiment and theory. Delayed neutron emission plays an important role in the determination of elemental abundances in the r-process and is also crucial in nuclear reactor technology. In addition, -delayed neutron emission from exotic nuclei is a powerful spectroscopic tool for the study of the daughter nucleus. In this work, decays of 46,47Cl were utilized to populate excited states in 45,46,47Ar, close to the neutron magic number , by means of both direct decay and -delayed neutron emission.
Studies on the breakdown of the conventional magic number below doubly-magic 48Ca (, ) have provided considerable insight into the evolution of nuclear structure toward the neutron dripline [5, 6]. At the center of the Island of Inversion [7], spectroscopic measurements have shown that 42Si (, ) has a low first energy [8, 9, 10]. Similarly, 44S (, ) has shown characteristics of a diminished gap, such as large collectivity [11, 12] and multiple shape coexistence [13, 14, 15].
Properties of the transitional nucleus 46Ar (, ) on the other hand, have proven to be more difficult to describe theoretically. Successful shell-model interactions in this region of the chart of nuclides overpredict the B strength by a factor of two [16, 17, 18] and this issue persists for 47Ar [19], 45Cl [20], and 43,44S [21, 12]. The data on excited states in 46Ar beyond the first are limited and come from 9Be(48Ca,46Ar)X [22], inverse-kinematics proton scattering [23], and 44Ar [24]. Although each study reported several -ray de-exictations, there is no consistency between the respective level schemes, beyond the transition and, tentatively, the transition. More information on the levels in 46Ar will provide critical benchmarks for shell model Hamiltonians aiming to describe the underlying mechanisms driving nuclear shell evolution from 48Ca into the Island of Inversion.
In the present work, the level structure of 46Ar was obtained from both the decay of 46Cl and the decay of 47Cl. From the decay of 46Cl, indications of the population of a previously-unreported negative-parity state in 46Ar was observed. New states were also observed in the decay of 47Cl, although no conjectures on their parities could be made. The decay of 47Cl to 47Ar was found to be very weak, with 98% of the total decay strength instead going toward delayed neutron emission. Despite the very small decay branch, we have tentatively identified two new -ray transitions in 47Ar. In addition, we observed transitions in 45Ar from 47Cl decay. These transitions were compared with those seen from 46Cl decay and differences are highlighted.
Shell model calculations were performed using the FSU [25] and [26] effective interactions with the valence space including the and shells. One particle was allowed to move from the shell to the shell to create opposite-parity states. For calculations using the interaction, both allowed Gamow-Teller (GT) and First Forbidden (FF) transitions were included when calculating -decay half-lives and delayed neutron emission probabilities. Inclusion of FF transitions in the calculations had the effect of reducing the calculated half-lives and neutron emission probabilities as they add a pathway to the decay.
II Experimental Setup
The experiment was carried out at the Coupled Cyclotron Facility at the National Superconducting Cyclotron Laboratory (NSCL) [27]. A 140-MeV/u 48Ca primary beam was fragmented on a thick Be target at the mid-acceptance position of the A1900 fragment separator [28]. The resulting secondary beam was purified using an achromatic Al wedge degrader at the intermediate dispersive image of the A1900 to select exotic isotopes of P, S, and Cl with neutron number around in two separate magnetic rigidity settings. Data from this experiment on the decays of 42,43,44P and 44,46S [29], 45Cl [30], and 43,45S [31] have been published and provide further details of the setup. Here, we report on the decays of 46,47Cl, which were produced in the higher rigidity setting with a momentum acceptance.
As shown in Fig. 1, particle identification was performed event-by-event using energy loss in two Si PIN detectors and their timing information relative to the plastic scintillator at the A1900 intermediate dispersive image. Downstream of the Si PIN detectors, the secondary beam particles, including the ions of interest, 46,47Cl, were implanted in the 40 strips x 40 strips pixelated Double-Sided Silicon Strip Detector (DSSD) belonging to the Beta Counting System (BCS) [32]. The implantation rate was about 150/s which allowed efficient correlation of implants and decays. An Al degrader was placed upstream to ensure that implants stopped near the middle of the 986-m thick DSSD. A Single-Sided Silicon Strip Detector (SSSD) was installed downstream to veto light particles punching through the DSSD. The BCS was surrounded by 16 HPGe Clover detectors to record -delayed rays. Energy and efficiency calibration was performed using standard -ray sources up to 3.5 MeV. The efficiency was about 5% at 1 MeV. The setup did not include neutron detection, and, hence, information about delayed neutron emission was obtained from delayed -ray transitions instead. The data were collected using the NSCL digital data acquisition system which provides timestamps for each channel, allowing coincidences and correlations to be built in the offline analysis [33].
III Experimental Results
III.1 Half-life measurements
Isotope | (ms) | ||||
Exp. | (keV) | GT+FF | GT | ||
46Cl | 216(16) ms | 0 | 271.4 | 345.9 | |
310 | 221.9 | 255.7 | |||
160 | 240.8 | 287.6 | |||
370 | 333.8 | 400.7 | |||
47Cl | 109.5(25) ms | 0 | 82.9 | 103.5 | |
30 | 109.5 | 129.3 |
The decay curves for 46,47Cl are shown in Figs. 2(a) and 2(b) respectively. These curves were generated by histogramming the time difference between the selected implanted ion (46Cl or 47Cl) and its decay product. The particle was required to be detected within the same pixel as the implant or in one of the neighboring eight pixels of the implant within a time correlation window of 2 s. Each decay curve was fitted with an expression to account for the exponential decay of the parent nucleus and the growth and decay of the daughter nuclei following the Bateman equations. The resulting half-life for 46Cl is 201(16) ms which agrees with the previous measurement of 223(37) ms [35, 36] within uncertainties and is slightly lower than the value of 232(2) ms from Ref. [37] but within 2. The half-life measured for 47Cl in this work is 109.5(25) ms, which is slightly higher than the literature value of 101(6) ms [37], but within their mutual uncertainties. In both cases, the exponential growth and decay of the daughter nuclei using the literature values for their half-lives [34] and an exponential background from long-lived activities were accounted for. Additionally, the 47Cl half-life was determined by gating on the 1553-keV -ray transition in the daughter, 46Ar (Fig. 2(c)). An exponential (parent decay) and a constant (background) were fitted to the curve, yielding a half-life of 109(5) ms, in excellent agreement with the ungated fit.
The calculated half-lives for 46Cl and 47Cl using the interaction [26] for different assumptions for the spin-parity of the parent ground state are listed in Table 1. The half-lives are calculated both with GT transitions alone and with GT+FF transitions and experimental values were used. For 46Cl, the calculated GT+FF values are closest to the experimental values of 223(37) ms [35, 36], and 232(2) ms [37], and 201(16) ms from this work if the ground state is either or . For 47Cl, the calculated value of 109.54 ms assuming a ground state and using GT+FF agrees remarkably well with the measured half-life in the present work of 109.5(25) ms. A ground state for 47Cl is also consistent with the systematics for the Cl isotopes [31].
III.2 delayed -ray transitions
-delayed -ray spectra correlated with 46,47Cl implants are shown in Figs. 3(a) and (b). Time correlation windows of 250 ms and 100 ms were used for 46Cl and 47Cl, respectively. These time intervals correspond to roughly one half-life for the two cases and highlight decays from the parent nucleus. For 46Cl decay (Fig. 3(a)), the spectrum is dominated by the 1553(2)-keV and 542(1)-keV transitions which are the ground-state decays of the first excited states in the daughter (46Ar) and the daughter (45Ar), respectively. For 47Cl decay (Fig. 3(b)), the spectrum is again dominated by the 1553(2)-keV transition in 46Ar ( daughter), signaling that the -delayed neutron emission channel is stronger than the decay to neutron-bound states. The 1231-keV transition represents the lowest one in 47Ar, the daughter.
- coincidences observed in 47Ar following 47Cl decay are shown in Fig. 4(a). A gate on the known 1231(1)-keV transition () in 47Ar highlights its coincidence with the 1602(2)-keV and 1935(2)-keV transitions. These transitions have not been reported previously, and correspond to new levels at 2833(2) keV and 3166(2) keV, respectively. The ground state of 47Cl is conjectured to be = from the present half-life measurement (see Table 1), and, thus, these states are candidates for positive-parity states populated by allowed Gamow-Teller (GT) transitions. Fig. 4(b) shows known - coincidences in the granddaughter nucleus 47K [38] which follows the decay of 47Ar confirming its production in the relatively weak branch of the decay of 47Cl.
The transition in 46Ar at 1553 keV is the most intense transition observed in both the 46Cl and 47Cl decays (Figs 3(a),(b)). The coincidences observed with the 1553-keV transition in the two cases are shown in Figs. 5(a) and (b). From 46Cl decay, a 1143(3)-keV transition was observed in coincidence with the 1553(2)-keV transition, as seen in Fig. 5(a). This could correspond to the 1140(20)-keV peak from Ref. [22]. 46Cl decay was also reported in a conference proceedings where 2006-keV and 3350-keV rays were observed in coincidence with each other and with the 1553-keV transition [39]. We were able to confirm these coincidence relationships (Fig. 5(b)) and further suggest that the 2009(3)-keV transition feeds into the 3347(3)-keV transition, although the reverse cannot be ruled out, as both transitions are equally intense within the estimated uncertainties. The statistics for 46Cl decay were rather low in this experiment and no further -ray transitions could be identified. The absolute feeding to the observed levels in 46Ar from 46Cl decay and the corresponding log values are given in Table 2.
(keV) | (I) | log |
1553(2) | 49(24)111should be considered an upper limit due to possible unobserved transitions | 8.34(33) [FF] |
2696(2) | 2.6(6) | 9.38() [FF] |
3249(3) | 3.7(8) | 9.10() [FF] |
4900(3) | 0 | |
6909(4) | 7.5(21) | 5.51() [GT] |
From 47Cl decay, several transitions were observed in coincidence with the 1553-keV peak in 46Ar, as seen in Fig. 5(c). Due to low statistics, these transitions were difficult to observe in an ungated spectrum. The 2133(2), 2311(2), 2517(2), and 2699(2)-keV transitions in the present work may correspond to the 2141(3), 2318(3), 2518(2), 2707(2)-keV transitions cited in Ref. [24], although they are somewhat lower in energy. The 2307(13) and 2692(16)-keV rays reported in the inverse kinematics proton scattering of 46Ar in Ref. [23] are similar in energy to the 2311 and 2699-keV transitions seen in the present work. The 3037(2), 3197(3), 3427(3), and 3627(3)-keV transitions are reported for the first time in this work. In Ref. [23] a 3430(26)-keV ray was observed in coincidence with the transition and could correspond to the 3427-keV transition seen here. However, other decays observed in Ref. [23] from that state were not observed in this work, leading us to suggest that they could be distinct. The relative intensities of the rays observed in coincidence with the 1553-keV transition are provided in Table 3 for both 46Cl and 47Cl decay.
47Cl decay | 46Cl decay | ||
(keV) | (keV) | ||
2133(2) | 52(16) | 1143(3) | 38(19) |
2311(2) | 40(14) | 1695(3) | 39(23) |
2517(2) | 99(23) | 2009(3) | 61(30) |
2699(2) | 100 | 3347(3) | 100 |
3037(2) | 86(23) | ||
3197(3) | 35(16) | ||
3427(3) | 30(15) | ||
3627(3) | 24(14) |
47Cl decay | 46Cl decay | ||
(keV) | (keV) | ||
1193(1) | 24(9) | 1193(1) | 22(13) |
1230(1) | 91(18) | 1230(1) | 100 |
1525(1) | 63(16) | 1525(1) | 27(16) |
2755(2) | 100 | 2755(2) | 63(31) |
799(1) | 68(20) |
The value for 47Cl is 16040(100) keV, while is 4.0(3) MeV [34]. This allowed us to observe -ray transitions in the daughter, 45Ar, which were also independently populated via the 46Cl channel ( = 7.96(10) MeV). For both cases, clean coincidences with the 542(1)-keV ray (de-excitation of the first excited state to the ground state in 45Ar) were observed. The 46Cl and 47Cl delayed -ray spectra gated on the 542(1)-keV transition are compared in Figs. 6(a) and (b), and the corresponding partial level schemes are also displayed. For 45Ar from 47Cl decay, only decays from positive-parity states above the 542-keV level were observed [30, 39]. For 45Ar populated from 46Cl decay, the rays from the same set of positive-parity states are seen along with a strong 799(1)-keV transition. Based on the 45Cl decay scheme reported in Refs. [30, 39], this ray is assigned to the de-excitation of the state at 1341(1) keV. Relative intensities of the -ray transitions observed in the 542-keV gate are provided in Table 4 for both cases.
IV Discussion
IV.1 46Ar
GT:46Cl () | GT:46Cl () | GT:46Cl () | GT:46Cl () | ||||||||
(keV) | (keV) | (keV) | (keV) | ||||||||
5570 | 5.37 | 5860 | 5.94 | 5570 | 5.64 | 5370 | 5.70 | ||||
6759 | 5.39 | 6902 | 5.63 | 5932 | 5.98 | - | - | - | |||
- | - | - | 5570 | 5.27 | 5116 | 5.86 | - | - | - | ||
- | - | - | 5932 | 5.81 | 6022 | 5.60 | - | - | - | ||
- | - | - | 6369 | 5.61 | 6369 | 5.48 | - | - | - | ||
- | - | - | (FF) | 1443 | 7.71 | (FF) | 1443 | 6.82 | (FF) | 1443 | 7.32 |
(FF) | 3496 | 6.89 | (FF) | 3496 | 8.62 | - | - | - | - | - | - |
The level scheme of 46Ar, which is the daughter of 46Cl and the daughter of 47Cl was established using - coincidences (Fig. 5) and energy and intensity balance. The partial level schemes with states in 46Ar populated in 46Cl decay (left) and 47Cl decay (right) are shown in Fig. 7. Interestingly, though not a complete surprise, the levels beyond the first observed from 46Cl decay and 47Cl decay are completely different. The low spin states predicted from shell model calculations using the FSU interaction [25] and the interaction [26] are displayed alongside for comparison.
Both interactions predict negative-parity levels in 46Ar occurring above an excitation energy of 4.5 MeV. The level at 6909(4) keV which is directly fed in the decay of 46Cl is a very likely candidate for a negative-parity state. The 46Cl ground state likely has = or , as was discussed in section III A. The intensity observed (Table 2) for this level corresponds to a log value of 5.51 which is consistent with an allowed GT transition. Log values from shell model calculations using the interaction are listed in Table 5 assuming different values for the 46Cl ground state. The decay from a ground state to a level at 6902 keV has a log value of 5.63 and to a state at 6369 keV has a log value of 5.61, making both possible candidates. The state at 6369 keV can also be populated from the decay of a ground state in 46Cl and is predicted to have a log of 5.48. Since the 4900-keV level is fed by this probable negative-parity state and decays to the level at 1553 keV, it will likely have a spin of 1 or 2 with either parity possible.
In Ref. [22] the energy was reported as 1570(5) keV, while the energy of the is quoted as 1558(9) keV in Ref. [23], 1554(1) keV in Ref. [24], and 1553(2) keV in this work. The state is tentatively placed at 3892(9) keV in Ref. [22], at 3866(16) keV in Ref. [23], and at 3872(3) keV in Ref. [24]. The transition analogous in energy in the current work (2311 keV) would place the state at 3864 keV. The shell model calculations place the at 3306 keV (FSU) or 3741 keV () in reasonable agreement.
The first excited level was tentatively placed at 2710(21) keV through comparison with level energies from shell model calculations in Ref. [22]. On the other hand, based on angular distribution analysis in the study of 44Ar [24], the level at 3695(3) keV was proposed as the state instead, while the state from Ref. [22] was not seen. In the current study, levels with similar energies to both the prior candidates, 2696 keV and 3686 keV, are observed in 46Cl decay and 47Cl decay, respectively. Shell model calculations generate several yrare states with both interactions agreeing on being the third excited state, although differing in energy. The (2948 keV) and (3623 keV) predictions from the FSU interaction agree well with the two experimental states in question. In our previous paper on of 45 Cl [30], we had identified the in 44Ar produced again in the channel at 2978 keV, while the FSU interaction predicted the state at 2717 keV, a difference of 261 keV. Similarly assigning the 2696 keV as in 46Ar compared to the calculated 2948 keV would be an energy difference of 252 keV making it a very plausible scenario. Of course, more experimentation is needed to definitively clarify the position of the excited states, which will impact our understanding of shell structure in this region and can be correlated to nuclear size and deformation.
Based on the experimental information we have, it is not feasible to comment on the spin or parity of other states populated in the 47Cl decay. With decays to the state, their spins can be constrained to be within 0 and 4 if they have positive parity and likely within 1 and 3 if they have negative parity. Based on relative intensities (see Table 3), it appears that the levels at 4070 keV, 4252 keV, and 4590 keV are favored over the others in the neutron emission process, though a definitive determination of the spin-parity cannot be made. In Ref. [24], these states were populated in a reaction on 44Ar and were thought to have positive parity, although no spin assignments were made. A positive parity is consistent with their population in the process, as will be discussed in Section IV.3.
IV.2 47Ar
The - coincidences observed in 47Ar following 47Cl decay were shown in Fig. 4(a). Based on these coincidences, a very tentative level scheme for 47Ar is proposed and shown in Fig. 8. Selected states predicted by shell-model calculations using the FSU interaction [25] and the interaction [26] are displayed alongside for comparison. The neutron separation energy in 47Ar is 3667(3) keV and, therefore, Gamow-Teller decays from the likely ground state of 47Cl would feed levels predicted to be close to or above the neutron separation energy. For example, the log value to the shell-model state at 3499 keV () is 5.66 and the log value to the state at 4202 keV is 5.77. These predicted positive-parity states are higher in energy than the experimentally observed states. Alternatively, the observed levels may have been populated by First Forbidden (FF) decay. The log values to the shell-model states with of (2281 keV), (2550 keV), (3479 keV) calculated using are 6.79, 6.55, and 6.86 respectively, making all of these predicted states possible candidates for the observed levels.
A state was reported at 1745 keV in a prior in-beam -ray spectroscopy experiment following one-proton knockout [40]. The same experiment proposed a state at 2186 keV and states at 1231 keV, 2763 keV, and 3437 keV [40]. These states are indicated by green lines in Fig. 8 and as can be seen they align closely in energy to levels with the same predicted by and FSU interactions. The previously-observed level at 1745 keV nicely corroborates with the shell-model state at 1371 keV from the interaction and 1890 keV state as predicted using the FSU interaction. The interaction provides candidate for the 2188 keV and 2763 keV states in the and states. This would leave open the possibility of the present 2833-keV and 3166-keV levels tentatively being the (2281 KeV) and (2550 keV) states though somewhat lower in energy. The equivalent states calculated using the FSU interaction are at 2365 keV () and 3074 keV ().
IV.3 -delayed neutron emission
In prior studies, the reported experimental probabilities for -delayed neutron emission showed dramatic changes as a function of neutron number: 24(4)% for 45Cl, 60(9)% for 46Cl, and then, very surprisingly, less than 3% for 47Cl [35]. From the present work, following the activity in the granddaughter nuclei from the three branches, , , and , we have estimated the values to be 75(15)% for 46Cl and 98(12)% for 47Cl. This smooth and increasing trend is what one would expect as neutron number increases. The measured values are compared with those from shell model calculations () with and without the inclusion of FF transitions in Table 6. Although the calculations reproduce the trend of increasing with neutron number, it is interesting to note that the values are underestimated in both cases. For 47Cl, the value becomes much smaller than the experimental value with the inclusion of FF transitions, suggesting that the FF contribution is overestimated in the calculation. Revisiting Fig. 3(b) emphasizes that FF transitions are weak for 47Cl decay, as the spectrum is dominated by rays from the daughter. In contrast, the measured values were better reproduced when FF transitions were included. This conflict needs to be explored further with higher-precision measurements of the neutron emission probabilities and the investigation of even more neutron-rich nuclei.
Isotope | (MeV) | (Exp.) | () | () | |
% | GT+FF, | GT, | |||
75(15) | 34.90 | 43.34 | |||
46Cl | 23.83 | 26.01 | |||
23.17 | 27.57 | ||||
32.28 | 38.28 | ||||
98(12) | 86.13 | 96.22 | |||
47Cl | 77.56 | 87.04 | |||
As noted before, a different set of excited states in 46Ar were populated in the decay of 47Cl than those populated in the direct decay of 46Cl. Comparison with prior studies and predictions from shell model calculations seem to favor positive parities for the states seen in decay. In Fig. 9, a simple illustration of how -delayed neutron emission can proceed based on the occupancies suggested by shell model calculations performed using the FSU interaction is shown.
The ground-state configuration of 47Cl () is with a filled neutron shell and a hole in the proton orbital. The states with proton excitations into the shell are energetically too high in 46Ar to be reached by GT decay (i.e conversion of one of the neutrons into a proton). Instead, a GT transformation changes one of the neutrons in either the or orbitals to a proton, filling the orbital, as shown by the arrow in the top cartoon. This leaves a hole in the or orbital resulting in a positive parity state of 47Ar. The ground state configuration of 47Ar () is with a full neutron shell and an unpaired neutron in the orbital. We can assume that the (log = 5.3) and (log = 5.5) states in 47Ar predicted at 4.95 MeV and 5.18 MeV, respectively, are populated in this allowed GT decay from the ground state of 47Cl. With equal to 3.667(3) MeV, these states lie above the neutron separation threshold in 47Ar.
In the next step, neutron emission from these states populates low-lying positive-parity states in 46Ar, as seen experimentally. For that, an unpaired neutron is emitted from the shell creating a positive parity state in the daughter nucleus 46Ar, as seen in the bottom cartoon in Fig. 9. This emission is energetically favored due to nuclear pairing and the low orbitals involved. Breaking a pair would typically cost an additional one or two MeV. According to shell model calculations using the FSU interaction, the state in 46Ar is a state with a full neutron shell (). Hence, the neutron emission has to be facilitated by a small admixture of configuration if the state is populated directly. Therefore, delayed neutron emission is probing and manifesting mixing between configurations ( and in this case). In a recent publication [41], IBM calculations including configuration mixing (IBM-CM) predicted low-lying states in 46Ar. In that work, it was shown that the yrast and states have a low admixture of intruder [n+2] space while the , , , and states are dominated by intruder configurations. This can explain the experimental observation of their population in the -delayed neutron emission channel.
A schematic view of -delayed neutron emission from 46Cl, which has possible ground-state spin-parities of and , is shown in Fig. 10. GT allowed decay would transform one of the neutrons into an proton, while transformation of the unpaired neutron in the orbital to a proton would be a FF transition. Following GT decay, the emission of a neutron would create a state in 45Ar due to the unpaired neutron. On the other hand, if the neutron is emitted, a state would be populated in the daughter nucleus due to the unpaired neutron. This simple picture seems to be consistent with the observed population of both positive-parity (e.g. 3297 keV, ) and negative-parity (e.g. 1341 keV, ) states, as shown in Fig. 6. The occupancy of the state from the shell-model calculation shows contributions from a hole along with an hole [30].
V Summary
The Beta Counting System surrounded by an array of 16 HPGe Clover detectors was used to study the and -delayed neutron decays of the exotic 46,47Cl isotopes at the National Superconducting Cyclotron Laboratory. The level structures of 45,46,47Ar were elucidated from the various decays, and comparisons were made to shell model calculations when possible. Following the activity in the granddaughter nuclei, the values for 46Cl and 47Cl are estimated to be 75(15)% and 98(12)%, respectively. The large for 47Cl is at odds with an extremely low prior measurement of . However the increasing trend for going from 45Cl to 47Cl, is expected as the neutron dripline is approached. The half-lives of 46,47Cl were measured and compared with calculations using the interaction for a variety of possible ground-state spins. When including both GT and FF transitions, good agreement between experiment and theory was found assuming spin-parities of or for the 46Cl ground state and a spin-parity of for the 47Cl ground state. These spin-parities values are also consistent with the observed decays. Two new -ray transitions in 47Ar were identified following the weak decay of 47Cl. It was also observed that the levels populated in 46Ar from 46Cl and 47Cl decay are completely different beyond the first excited state. Differences were also seen in the states populated in 45Ar from 46Cl decay compared to 47Cl decay. Attempts were made to explain states populated in delayed neutron emission using a simple shell-model picture.
The large values observed for 46,47Cl indicate a very small decay branch to bound states, likely through FF transitions. This observation seems to be consistent for 47Cl decay. However, for 46Cl, a large branch to the low-lying state was observed, somewhat contradictory to the large value. Furthermore, shell model calculations using the interaction which reproduced the measured half-lives after including FF decays severely underestimated the measured values. This highlights that further work is necessary, both in experiments and in the refinement of theoretical approaches, to quantify the impact of FF transitions. This region of neutron-rich isotopes around the Island of Inversion remains a fertile testing ground for state-of-the-art nuclear theory. With the start of operations at the new Facility for Rare Isotope Beams (FRIB), nuclei with large -delayed neutron branches, like 46Cl and 47Cl isotopes studied in the present work, will feature prominently in experimental campaigns. Key experimental devices like the FRIB Decay Station Initiator will allow measurements of both neutrons and rays from these exotic systems, allowing for complete decay spectroscopy [42]. With the new radioactive-beam facilities beginning to provide much-needed experimental data, a clearer picture of -delayed neutron decay and the role of FF transitions may soon emerge.
Acknowledgments
We thank the NSCL operations team and the A1900 team, especially Tom Ginter for the production and optimization of the secondary beam. This work was supported by the US National Science Foundation under Grant Nos. PHY-2012522 (FSU) and PHY-1848177 (CAREER), the US Department of Energy, Office of Science, Office of Nuclear Physics under award Nos. DE-SC0020451 (FRIB), DE-FG02-94ER40848 (UML), DE-AC52-07NA27344 (LLNL), DE-AC02-06CH11357 (ANL), and DE-SC0009883 (FSU) and also by the US Department of Energy National Nuclear Security Administration Grant No. DOE-DE-NA0003906 and the Nuclear Science and Security Consortium under Award No. DE-NA0003180.